OpenStudy (anonymous):

what is the solution to 2x-5y=24 and 15x-17y=16

OpenStudy (texaschic101):

elimination method... 2x - 5y = 24 ---> (15)2x - 5y = 24 15x - 17y = 16 --> (-2)15x - 17y = 16 ---------------- 30x - 75y = 360 (result of multiplying by 15) -30x + 34y = - 32 (result of multiplying by -2) ----------------add 0 - 41y = 328 -41y = 328 (divide by -41 on both sides to get y by itself) -41y/-41 = 328/-41 y = -8 now sub -8 in for y in either of the original equations... 2x - 5y = 24 2x - 5(-8) = 24 2x + 40 = 24 2x + 40 - 40 = 24 - 40 2x = 24 - 40 2x = - 16 2x/2 = -16/2 x = - 8 check... 2x - 5y = 24 2(-8) - 5(-8) = 24 -16 + 40 = 24 24 = 24 (correct) SOLUTION : (-8,-8) ---> x = -8 and y = -8